# On the Global Optimality of Linear Policies for Sinkhorn Distributionally Robust Linear Quadratic Control

### v1

@@ Abstract
The Linear Quadratic Gaussian (LQG) regulator is a cornerstone of optimal control theory, yet its performance can degrade significantly when the noise distributions deviate from the assumed Gaussian model. To address this limitation, this work proposes a distributionally robust generalization of the finite-horizon LQG control problem. Specifically, we assume that the noise distributions are unknown and belong to ambiguity sets defined in terms of an entropy-regularized Wasserstein distance centered at a nominal Gaussian distribution. By deriving novel bounds on this Sinkhorn discrepancy and proving structural and topological properties of the resulting ambiguity sets, we establish global optimality of linear policies. Numerical experiments showcase improved distributional robustness of our control policy.
@@ I Introduction
In the presence of model misspecifications, however, the LQG solution can be extremely fragile [2]. Classical $\mathcal{H}_{\infty}$ control [3] addresses this concern by shifting from a stochastic to an adversarial uncertainty model and minimizing the worst-case cost across bounded-energy disturbances. While provably robust, $\mathcal{H}_{\infty}$ methods tend to be overly conservative in practice as they optimize for least favorable uncertainty realization. Motivated by this observation, several approaches have been proposed to balance nominal performance and robustness, including mixed $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ formulations [4, 5], risk-sensitive control [6], and regret minimization methods [7, 8, 9].
@@
Among these approaches, recent work on distributionally robust (DR) control promises to combine the advantages of stochastic and adversarial uncertainty models by robustifying in the probability space. To achieve this, this paradigm considers the problem of minimizing the expected cost under the most averse distribution within a given ambiguity set—a set of distributions that are sufficiently close, in appropriate sense, to a nominal one. For instance, [10] studied DR control of constrained stochastic systems with ambiguity sets comprising all distributions sharing the same first two moments. To account for full distributional information, the works [11, 12, 13] instead employ ambiguity sets defined in terms of $f$-divergence, whereas [14, 15, 16, 17, 18, 19] rely on optimal transport metrics in light of their proven expressiveness and out-of-distribution guarantees.
@@
Inspired by these results, we study a generalization of the finite-horizon LQG control problem, where the noise distributions are unknown and belong to Sinkhorn ambiguity sets [20, 21] centered at nominal Gaussian distributions. Our main contribution is to establish global optimality of linear policies for this Sinkhorn DR LQG control problem, generalizing the results of [14] to the case where the definition of Wasserstein distance includes a Kullback-Leibler (KL) regularization term [22]. Towards deriving our main result, we first construct a lower bound to our DR LQG problem leveraging results from regularized optimal transport [23]. Second, we prove a Gelbrich-type inequality that bounds from below the Sinkhorn discrepancy between two probability distributions when one of them is Gaussian. This allows us to obtain an upper bound to our DR LQG problem. Third, we show convexity and compactness of the resulting entropy-regularized Gelbrich ambiguity set. These properties allow us to conclude, using a “sandwich” argument similar to [14], that the Sinkhorn DR LQG admits a Nash equilibrium in the class of linear policies.
@@
Notation: Throughout the paper, we denote the set of (zero-mean) probability distributions supported on a measurable set $\mathcal{Z}$ by $\mathcal{P}(\mathcal{Z})$ (resp. $\mathcal{P}_{0}(\mathcal{Z})$). We write $\mu\ll\nu$ to denote that a measure $\mu$ is absolutely continuous with respect to $\nu$. The convolution product between two probability measures is represented by $\mu*\nu$. For $n\in\mathbb{N}$, we write $[n]$ to denote the set of indices $\{0,\dots,n-1\}$. The space of all $d\times d$ positive (semi)definite matrices is denoted by $\mathbb{S}_{+}^{d}$ (resp. $\mathbb{S}^{d}$). We denote by $\|\cdot\|$ the Euclidean norm. The determinant of a square matrix $A$ is denoted by $|A|$. Given $A\in\mathbb{S}^{d}$, we denote by $\{\lambda_{i}(A)\}_{i=1}^{d}$ its eigenvalues and let $\lambda_{\text{max}}(A)=\max_{i}\lambda_{i}(A)$.
@@ II Preliminaries
As noted in [20, Remark 2], any choice of $\mathbb{P}\ll\mu$ in (1) is equivalent up to a constant. Hence, as in our DR LQG problem the center distribution is assumed to be known, we let $\mu=\mathbb{P}$ without loss of generality. Aligned with [20, 21], in the following we further assume $\nu\sim\mathcal{N}(0,\Sigma)$; this choice ensures that $W_{\epsilon}(\mathbb{P},\mathbb{Q})$ is finite for any $\mathbb{Q}$ in the ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$, independently of its support.
@@
$\mathrm{KL}(\gamma\|\mu\times\nu)=\mathrm{KL}(\gamma\|\mathbb{P}\times\mathbb{Q})-\mathrm{KL}(\mathbb{P}\times\mathbb{Q}\|\mu\times\nu)\,;$
@@
Let $\nu\sim\mathcal{N}(0,\Sigma)$ where $\Sigma\in\mathbb{S}^{d}_{+}$. The entropy-regularized Gelbrich divergence between two probabilities $\mathbb{P}_{1},\mathbb{P}_{2}\in\mathcal{P}_{0}(\mathbb{R}^{d})$ with covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}_{+}^{d}$ is
@@ III Problem Formulation
For ease of presentation, we collect all exogenous random vectors in the variable $\delta=(x_{0},w_{0},\dots,w_{T-1},v_{0},\dots,v_{T-1})$. We assume that all entries of $\delta$ are mutually independent. Differently from classical LQG theory, however, we assume that their true distributions are unknown and belong to a Sinkhorn ambiguity set $\mathcal{S}$ centered at a nominal distribution $\hat{\mathbb{P}}=\hat{\mathbb{P}}_{x_{0}}\otimes\left(\otimes_{t=0}^{T-1}\hat{\mathbb{P}}_{w_{t}}\right)\otimes\left(\otimes_{t=0}^{T-1}\hat{\mathbb{P}}_{v_{t}}\right)$, with $\hat{\mathbb{P}}_{x_{0}}=\mathcal{N}(0,\hat{X}_{0})$, $\hat{\mathbb{P}}_{w_{t}}=\mathcal{N}(0,\hat{W}_{t})$, and $\hat{\mathbb{P}}_{v_{t}}=\mathcal{N}(0,\hat{V}_{t})$, for all $t\in[T]$. [^1] Following [14, 15, 13, 12], we consider only distributions with zero mean; our results extend to the non-zero mean case with minor modifications. Specifically, for user-defined radii $\rho_{x_{0}}\geq\underline{\rho}_{x_{0}},\rho_{w_{t}}\geq\underline{\rho}_{w_{t}}\rho_{v_{t}}\geq\underline{\rho}_{v_{t}}$ and regularization parameter $\epsilon\geq 0$, we define the ambiguity set $\mathcal{S}$ as $\mathcal{S}_{x_{0}}\otimes\left(\otimes_{t=0}^{T-1}\mathcal{S}_{w_{t}}\right)\otimes\left(\otimes_{t=0}^{T-1}\mathcal{S}_{v_{t}}\right)$ where
@@
where $Q_{t},Q_{T}\in\mathbb{S}^{d}$ and $R_{t}\in\mathbb{S}^{m}_{+}$ represent state and input weight matrices, respectively. With these definitions in place, we formulate the Sinkhorn DR LQG control problem as
@@
where $\mathcal{U}_{y}$ denotes the set of all of feasible control inputs $\bm{u}=(u_{0},\dots,u_{T-1})$. In particular, (4) can be interpreted as a zero-sum game between the control designer, who selects a causal policy to minimize the expected cost, and an adversary, who chooses the noise distributions in $\mathcal{S}$ that maximize such cost.
@@ IV Analysis of the Sinkhorn DR LQG problem
We now present our main results. We first re-parametrize (4) in terms of the purified observations [27]. Then, we construct two auxiliary problems over the class of linear policies that provide lower and upper bounds to (4). Last, using a “sandwich” argument, we show that the optimal value of these two auxiliary problems coincide, implying that (4) admits a Nash equilibrium and that linear policies are globally optimal.
@@ IV-A Purified output re-parametrization
To ease analysis of (4), we rewrite $u_{t}$ in terms of the purified observations instead of the actual observations $y_{t:0}$. To define these new variables, we introduce a noise-free copy of (3) as
@@
with state $\hat{x}_{t}\in\mathbb{R}^{d}$ and output $\hat{y}_{t}\in\mathbb{R}^{p}$, initialized at $\hat{x}_{0}=0$ and with the same input $u_{t}$ as the original system. We then define the purified output $\eta_{t}$ at time $t$ as $\eta_{t}=y_{t}-\hat{y}_{t}$, and let $\bm{\eta}=(\eta_{0},\dots,\eta_{T-1})$. This representation proves useful because, as shown in [28, Proposition II.1], every measurable function of $y_{0},\dots,y_{t}$ can be equivalently expressed as a measurable function of $\eta_{0},\dots,\eta_{t}$, and vice versa—yet differently from $\mathbf{y}=(y_{0},\dots,y_{T-1})$, $\bm{\eta}$ is independent of the inputs. In fact, using (3) and (5) recursively, one can show that $\bm{\eta}$ only depends on the exogenous vectors $\mathbf{w}=(x_{0},w_{0},\dots,w_{T-1})$ and $\mathbf{v}=(v_{0},\dots,v_{T-1})$. In particular, it holds that $\bf\bm{\eta}=Dw+v$, where $\bf D=CG$ and $\bf C$ and $\bf G$ are matrices defined in the Appendix.
@@
Classical min-max inequality states that $d^{\star}\leq p^{\star}$. In the next sections, we prove that such relation actually holds with equality—despite the fact that $\mathcal{U}_{\eta}$ is an infinite-dimensional function space and $\mathcal{S}$ is an infinite-dimensional set of non-parametric probability distributions. In particular, our analysis reveals that that there exists a Nash equilibrium of the zero-sum game (4) in the form $(\mathbf{u}^{\star},\mathbb{P}^{\star})$, where $\bf u^{\star}=U^{\star}\bm{\eta}+q^{\star}$ is an affine policy of $\bm{\eta}$ and $\mathbb{P}^{\star}$ is Gaussian. To do so, we first construct a lower-bound to (7) and an upper-bound to (6), and then show that their optimal values coincide.
@@ IV-B Construction of a lower bound for $d^{\star}$
Compared to (7), in (8) we restricted the feasible set in the outer maximization. Hence, we have $\underline{d}^{\star}\leq d^{\star}$. On the other hand, (8) is still an infinite-dimensional optimization problem. In the remaining of this section, we show that this problem can be reformulated as a finite-dimensional one by exploiting known closed-form expressions for the Sinkhorn discrepancy between two normal distributions.
@@
  •   Proof:
    As observed in Remark 1, regularizing (1) with the negative differential entropy or a KL term leads to the same optimal transport plan. Hence, the expression for $\gamma_{0}$ follows from [23, Theorem 2.2]. Substituting this optimal coupling in the definition of Sinkhorn discrepancy in (1), we obtain
@@
    By inspection, combining the expressions above as per (1) yields (2), which concludes the proof. ∎
Let us define the matrices $\mathbf{M}=\mathbf{R}+\mathbf{H}^{\top}\mathbf{QH}\in\mathbb{R}^{mT}$, $\mathbf{F}_{1}=\mathbf{D}^{\top}\mathbf{U}^{\top}\mathbf{R}\mathbf{U}\mathbf{D}+(\mathbf{HUD}+\mathbf{G})^{\top}\!\mathbf{Q}(\mathbf{HUD}+\mathbf{G})\in\mathbb{S}^{d(T+1)}$, and $\mathbf{F}_{2}=\mathbf{U}^{\top}\mathbf{R}\mathbf{U}+\mathbf{U}^{\top}\mathbf{H}^{\top}\mathbf{QHU}\in\mathbb{S}^{pT}$ for brevity. Moreover, we denote the set of causal feedback matrices with $\mathcal{U}^{\text{lin}}$. With this notation in place, we are now ready to reformulate (8) as a finite-dimensional optimization problem.
@@
$\mathcal{G}_{W}$ $=\big{\{}\mathbf{W}\in\mathbb{S}_{+}^{d(T+1)}:\ \mathbf{W}=\operatorname{diag}(X_{0},W_{0},\ldots,W_{T{-}1}),$
$\quad\quad X_{0}\in\mathbb{S}_{+}^{d},W_{t}\in\mathbb{S}_{+}^{d},\ G_{\epsilon}(\hat{X}_{0},X_{0})\leq\rho_{x_{0}},$
@@
$\mathcal{G}_{V}$ $=\big{\{}\mathbf{V}\in\mathbb{S}_{+}^{pT}:\ \mathbf{V}=\operatorname{diag}(V_{0},\ldots,V_{T{-}1}),\ V_{t}\in\mathbb{S}_{+}^{p},$
@@
    Following the same argument as [14, Proposition 3.2], the proof is concluded by rewriting the expectation of a quadratic form as a trace and by replacing the ambiguity set $\mathcal{G}$ by (10) and (11). ∎

### v2

@@ Abstract
The Linear Quadratic Gaussian (LQG) regulator is a cornerstone of optimal control theory, yet its performance can degrade significantly when the noise distributions deviate from the assumed Gaussian model. To address this limitation, this work proposes a distributionally robust generalization of the finite-horizon LQG control problem. Specifically, we assume that the noise distributions are unknown and belong to ambiguity sets defined in terms of a Sinkhorn discrepancy centered at a nominal Gaussian distribution. By deriving novel bounds on this entropy-regularized Wasserstein distance and proving structural and topological properties of the resulting ambiguity sets, we establish global optimality of linear policies for Sinkhorn distributionally robust LQG. Numerical experiments showcase improved distributional robustness of our control policy.
@@ I Introduction
In the presence of model misspecifications, however, the LQG solution can be extremely fragile [9]. Classical $\mathcal{H}_{\infty}$ control [32] addresses this concern by shifting from a stochastic to an adversarial uncertainty model and minimizing the worst-case cost across bounded-energy disturbances. While provably robust, $\mathcal{H}_{\infty}$ methods tend to be overly conservative in practice as they optimize for the least favorable uncertainty realization. Motivated by this observation, several approaches have been proposed to balance nominal performance and robustness, including mixed $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ formulations [3, 8], risk-sensitive control [17], and regret minimization methods [24, 15, 23].
@@
Among these approaches, recent work on distributionally robust (DR) control promises to combine the advantages of stochastic and adversarial uncertainty models by robustifying in the probability space. To achieve this, the DR control paradigm considers the problem of minimizing the expected cost under the most averse distribution within a given ambiguity set—a set of distributions that are sufficiently close, in an appropriate sense, to a nominal one. For instance, [29] studied DR control of constrained stochastic systems with ambiguity sets comprising all distributions sharing the same first two moments. To account for full distributional information, the works [25, 10, 11] instead employ ambiguity sets defined in terms of $f$-divergence, whereas [28, 21, 31, 5, 1, 19] rely on optimal transport metrics in light of their proven expressiveness and out-of-distribution guarantees.
@@
Inspired by these results and motivated by the recent application of the Sinkhorn discrepancy for DR learning [30, 14] and control [6], we study a generalization of the finite-horizon LQG control problem, where the noise distributions are unknown and belong to Sinkhorn ambiguity sets centered at nominal Gaussian distributions. Our main contribution is to establish the global optimality of linear policies for the Sinkhorn DR LQG problem. While our proof follows a “sandwich” argument akin to [28], the extension from Wasserstein DR LQG to Sinkhorn DR LQG is far from direct: the introduction of the KL regularization term fundamentally alters the geometry of the ambiguity set [^1] For instance, it is well-known that the worst-case distribution in Wasserstein DR optimization is finitely supported when so is the center distribution [12], while the Sinkhorn ambiguity set with a Gaussian reference measure only contains continuous distributions. and invalidates several key inequalities used in [28]. To recover a comparable optimality result, we develop new tools specific to the Sinkhorn framework. In particular, we first construct a novel lower bound to our DR LQG problem leveraging results from regularized optimal transport [7]; second, we establish a Gelbrich-type inequality for the Sinkhorn discrepancy, which is not available in the literature and is essential for deriving an upper bound to our DR LQG problem; last, we prove convexity and compactness of the resulting entropy-regularized Gelbrich ambiguity set. Taken together, these results allow us to conclude that the Sinkhorn DR LQG admits a Nash equilibrium in the class of linear policies.
@@
Notation: Throughout the paper, we denote the set of (zero-mean) probability distributions supported on a measurable set $\mathcal{Z}$ by $\mathcal{P}(\mathcal{Z})$ (resp. $\mathcal{P}_{0}(\mathcal{Z})$). We write $\mu\ll\nu$ to denote that a measure $\mu$ is absolutely continuous with respect to $\nu$. The convolution product between two probability measures is represented by $\mu*\nu$. For $n\in\mathbb{N}$, we write $[n]$ to denote the set of indices $\{0,\dots,n-1\}$. The space of all $d\times d$ positive (semi)definite matrices is denoted by $\mathbb{S}_{++}^{d}$ (resp. $\mathbb{S}^{d}_{+}$). We denote by $\|\cdot\|$ the Euclidean norm. The determinant of a square matrix $A$ is denoted by $|A|$. Given $A\in\mathbb{S}^{d}$, we denote by $\{\lambda_{i}(A)\}_{i=1}^{d}$ its eigenvalues and let $\lambda_{\text{max}}(A)=\max_{i}\lambda_{i}(A)$.
@@ II Preliminaries
As noted in [30, Remark 2], any choice of $\mathbb{P}\ll\mu$ in (1) is equivalent up to a constant. Hence, as in our DR LQG problem the center distribution is assumed to be known, we let $\mu=\mathbb{P}$ without loss of generality. Aligned with [30, 6], in the following we further assume $\nu\sim\mathcal{N}(0,\Sigma)$; this choice ensures that $W_{\epsilon}(\mathbb{P},\mathbb{Q})$ is finite for any $\mathbb{Q}$ in the ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$, independently of its support.
@@
$\mathrm{KL}(\gamma\|\mu\times\nu)=\mathrm{KL}(\gamma\|\mathbb{P}\times\mathbb{Q})+\mathrm{KL}(\mathbb{P}\times\mathbb{Q}\|\mu\times\nu)\,;$
@@
Let $\nu\sim\mathcal{N}(0,\Sigma)$ where $\Sigma\in\mathbb{S}^{d}_{++}$. The entropy-regularized Gelbrich divergence between two probabilities $\mathbb{P}_{1},\mathbb{P}_{2}\in\mathcal{P}_{0}(\mathbb{R}^{d})$ with covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}_{++}^{d}$ is
@@
$\underline{\rho}=\tfrac{\epsilon}{2}\big(\operatorname{Tr}\!\big(\Sigma^{-1}(\hat{\Sigma}+\tfrac{\epsilon}{2}I)\big)-d+\log|\Sigma|-d\log(\tfrac{\epsilon}{2})\big)\,.$ (3)
This value represents the smallest radius $\rho$ such that the ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$ is non-empty. Last, we remark that, when $\epsilon\rightarrow 0$, Definition 3 is equivalent to the squared Gelbrich distance, see [13, Theorem 2.1].
@@ III Problem Formulation
For ease of presentation, we collect all exogenous random vectors in the variable $\delta=(x_{0},w_{0},\dots,w_{T-1},v_{0},\dots,v_{T-1})$. We assume that all entries of $\delta$ are mutually independent. Differently from classical LQG theory, however, we assume that their true distributions are unknown and belong to a Sinkhorn ambiguity set $\mathcal{S}$ centered at a nominal distribution $\hat{\mathbb{P}}=\hat{\mathbb{P}}_{x_{0}}\otimes\left(\otimes_{t=0}^{T-1}\hat{\mathbb{P}}_{w_{t}}\right)\otimes\left(\otimes_{t=0}^{T-1}\hat{\mathbb{P}}_{v_{t}}\right)$, with $\hat{\mathbb{P}}_{x_{0}}=\mathcal{N}(0,\hat{X}_{0})$, $\hat{\mathbb{P}}_{w_{t}}=\mathcal{N}(0,\hat{W}_{t})$, and $\hat{\mathbb{P}}_{v_{t}}=\mathcal{N}(0,\hat{V}_{t})$, for all $t\in[T]$. [^2] Following [28, 21, 11, 10], we consider only distributions with zero mean; our results extend to the non-zero mean case with minor modifications. Specifically, for any regularization parameter $\epsilon\geq 0$ and user-defined radii $\rho_{x_{0}}\geq\underline{\rho}_{x_{0}},\rho_{w_{t}}\geq\underline{\rho}_{w_{t}},\rho_{v_{t}}\geq\underline{\rho}_{v_{t}}$, where $\underline{\rho}_{x_{0}},\underline{\rho}_{w_{t}}$, and $\underline{\rho}_{v_{t}}$ are defined according to (3) with $\hat{X}_{0}$, $\hat{W}_{t}$, and $\hat{V}_{t}$ in place of $\hat{\Sigma}$, respectively, we define the ambiguity set $\mathcal{S}$ as $\mathcal{S}_{x_{0}}\otimes\left(\otimes_{t=0}^{T-1}\mathcal{S}_{w_{t}}\right)\otimes\left(\otimes_{t=0}^{T-1}\mathcal{S}_{v_{t}}\right)$ where
@@
where $Q_{t},Q_{T}\in\mathbb{S}^{d}_{+}$ and $R_{t}\in\mathbb{S}^{m}_{++}$ represent state and input weight matrices, respectively. With these definitions in place, we formulate the Sinkhorn DR LQG control problem as
@@
where $\mathcal{U}_{y}$ denotes the set of all feasible control inputs $\bm{u}=(u_{0},\dots,u_{T-1})$. In particular, (5) can be interpreted as a zero-sum game between the control designer, who selects a causal policy to minimize the expected cost, and an adversary, who chooses the noise distributions in $\mathcal{S}$ that maximize such cost.
@@ IV Analysis of the Sinkhorn DR LQG problem
We now present our main results. We first re-parametrize (5) in terms of the purified observations [2, Section 14.4.2]. Then, we construct two auxiliary problems over the class of linear policies that provide lower and upper bounds to (5). Last, using a “sandwich” argument, we show that the optimal value of these two auxiliary problems coincide, implying that (5) admits a Nash equilibrium and that linear policies are globally optimal.
@@ IV-A Purified output re-parametrization
To ease analysis of (5), we rewrite $u_{t}$ in terms of the purified observations instead of the actual observations $y_{0:t}$. To define these new variables, we introduce a noise-free copy of (4) as
@@
with state $\hat{x}_{t}\in\mathbb{R}^{d}$ and output $\hat{y}_{t}\in\mathbb{R}^{p}$, initialized at $\hat{x}_{0}=0$ and with the same input $u_{t}$ as the original system. We then define the purified output $\eta_{t}$ at time $t$ as $\eta_{t}=y_{t}-\hat{y}_{t}$, and let $\bm{\eta}=(\eta_{0},\dots,\eta_{T-1})$. This representation proves useful because, as shown in [16, Proposition II.1], every measurable function of $y_{0},\dots,y_{t}$ can be equivalently expressed as a measurable function of $\eta_{0},\dots,\eta_{t}$, and vice versa—yet differently from $\mathbf{y}=(y_{0},\dots,y_{T-1})$, $\bm{\eta}$ is independent of the inputs. In fact, using (4) and (6) recursively, one can show that $\bm{\eta}$ only depends on the exogenous vectors $\mathbf{w}=(x_{0},w_{0},\dots,w_{T-1})$ and $\mathbf{v}=(v_{0},\dots,v_{T-1})$. In particular, it holds that $\bf\bm{\eta}=Dw+v$, where $\bf D=CG$ and $\bf C$ and $\bf G$ are matrices defined in the Appendix.
@@
Classical min-max inequality states that $d^{\star}\leq p^{\star}$. In the next sections, we prove that such relation actually holds with equality—despite the fact that $\mathcal{U}_{\eta}$ is an infinite-dimensional function space and $\mathcal{S}$ is an infinite-dimensional set of non-parametric probability distributions. In particular, our analysis reveals that there exists a Nash equilibrium of the zero-sum game (5) in the form $(\mathbf{u}^{\star},\mathbb{P}^{\star})$, where $\bf u^{\star}=U^{\star}\bm{\eta}+q^{\star}$ is an affine policy of $\bm{\eta}$ and $\mathbb{P}^{\star}$ is Gaussian. To do so, we first construct a lower-bound to (8) and an upper-bound to (7), and then show that their optimal values coincide.
@@ IV-B Construction of a lower bound for $d^{\star}$
Compared to (8), in (9) we restricted the feasible set in the outer maximization. Hence, we have $\underline{d}^{\star}\leq d^{\star}$. On the other hand, (9) is still an infinite-dimensional optimization problem. In the remainder of this section, we show that this problem can be reformulated as a finite-dimensional one by exploiting known closed-form expressions for the Sinkhorn discrepancy between two normal distributions.
@@
As observed in Remark 1, regularizing (1) with the negative differential entropy or a KL term leads to the same optimal transport plan. Hence, the expression for $\gamma_{0}$ follows from [7, Theorem 2.2]. Substituting this optimal coupling in the definition of Sinkhorn discrepancy in (1), we obtain
@@
By inspection, combining the expressions above as per (1) yields (2), which concludes the proof. ∎ Let us define the matrices $\mathbf{M}=\mathbf{R}+\mathbf{H}^{\top}\mathbf{QH}\in\mathbb{R}^{mT}$, $\mathbf{F}_{1}=\mathbf{D}^{\top}\mathbf{U}^{\top}\mathbf{R}\mathbf{U}\mathbf{D}+(\mathbf{HUD}+\mathbf{G})^{\top}\!\mathbf{Q}(\mathbf{HUD}+\mathbf{G})\in\mathbb{S}_{+}^{d(T+1)}$, and $\mathbf{F}_{2}=\mathbf{U}^{\top}\mathbf{R}\mathbf{U}+\mathbf{U}^{\top}\mathbf{H}^{\top}\mathbf{QHU}\in\mathbb{S}_{+}^{pT}$ for brevity. Moreover, we denote the set of causal feedback matrices by $\mathcal{U}^{\text{lin}}$. With this notation in place, we are now ready to reformulate (9) as a finite-dimensional optimization problem.
@@
$\mathcal{G}_{W}$ $=\big\{\mathbf{W}\in\mathbb{S}_{++}^{d(T+1)}:\ \mathbf{W}=\operatorname{diag}(X_{0},W_{0},\ldots,W_{T{-}1}),$
$\quad\quad X_{0}\in\mathbb{S}_{++}^{d},W_{t}\in\mathbb{S}_{++}^{d},\ G_{\epsilon}(\hat{X}_{0},X_{0})\leq\rho_{x_{0}},$
@@
$\mathcal{G}_{V}$ $=\big\{\mathbf{V}\in\mathbb{S}_{++}^{pT}:\ \mathbf{V}=\operatorname{diag}(V_{0},\ldots,V_{T{-}1}),\ V_{t}\in\mathbb{S}_{++}^{p},$
@@
Following the same argument as [28, Proposition 3.2], the proof is concluded by rewriting the expectation of a quadratic form as a trace and by replacing the ambiguity set $\mathcal{G}$ by (11) and (12). ∎